# A Multi-Level Superoptimizer for Tensor Programs

### v2

@@ 3. Case Study: RMSNorm
Figure 5. An overview of the $\mu$Graph generator.
RMSNorm are generally followed by a matrix multiplication (MatMul). Figure 4a shows the computation graph of an RMSNorm and a MatMul operator, where $X$ is the input tensor, and $G$ and $W$ denote two weight tensors. Existing ML compilers generally launch two separate kernels for RMSNorm and MatMul computations, since both of them internally perform reduction across an input dimension, making it challenging to fuse their computations in a single kernel. This approach requires storing intermediate results (i.e., $Y$) in device memory as different kernels cannot share data in shared memory or register files.
@@
We highlight the key differences between the $\mu$Graph discovered by Mirage and the original $\mu$Graph. These differences involve discovering new custom kernels and combining algebraic and schedule transformations, making it infeasible to discover the final $\mu$Graph by separately considering algebraic and schedule transformations. First, Mirage reorders MatMul and the division of RMSNorm by leveraging the commutativity of matrix multiplication and element-wise division (algebraic transformation). Second, Mirage performs the accumulation in the root mean square (i.e., $\sum_{j}X_{j}^{2}$) and the accumulation in the matrix multiplication (i.e., $\sum_{i}X_{i}G_{i}W_{ik}$) in parallel (schedule transformation), avoiding writing the accumulation results to device memory. Next, Mirage instantiates a thread graph to perform a sequence of element-wise operators while maintaining all intermediate results in register files (schedule transformation). Finally, the best discovered $\mu$Graph uses a new custom kernel to fuse the computation of RMSNorm and MatMul, reducing device memory access and kernel launch overheads. This $\mu$Graph outperforms the hand-written kernels in existing systems by 1.9$\times$ and 1.6$\times$ on NVIDIA A100 and H100 GPUs.
4. Expression-Guided $\mu$Graph Generator
@@
3:$E_{O}\leftarrow\mathrm{AbstractExpression}(G_{\mathsf{ref}})$
4:$\mathcal{S}_{0},\mathcal{S}\leftarrow\varnothing,G\leftarrow\mathrm{Inputs}(G_{\mathsf{ref}})$
5:GenerateKernelOrBlockOps($\mathcal{S}_{0},G$)
6:ConstructThreadGraphs($\mathcal{S}_{0},\mathcal{S}$)
7:
8:function GenerateKernelOrBlockOps($\mathcal{S}_{0},G$)
9:  $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{G\}$
@@
12:    GenerateNextKernelOperator($G\cup\{o\}$)
@@
[…]
18:  return Invalid
19:function ConstructThreadGraphs($\mathcal{S}_{0},\mathcal{S}$)
20:  $\mathcal{P}\leftarrow$ pre-defined patterns
21:  for all $G\in\mathcal{S}_{0}$ do
22:   $G_{\mathsf{fused}}\leftarrow G$
23:   for all $(G_{i},O_{i})\in\mathcal{P}$ do
24:    for all subgraph $G^{\prime}$ of $G$ matching $G_{i}$ do
25:     Substitute $G^{\prime}$ with $O_{i}$ in $G_{\mathsf{fused}}$
26:   $S\leftarrow S\cup\{G_{\mathsf{fused}}\}$
This section introduces the Mirage $\mu$Graph generator, which automatically discovers potential $\mu$Graphs for an input tensor program. To generate $\mu$Graphs that capture optimizations at all of the kernel, block, and thread levels, Mirage must explore a significantly larger search space than existing superoptimizers that only consider optimizations at the kernel level. Mirage employs two key techniques to address this challenge. First, based on an important observation that optimizations at the kernel and block levels are substantially more critical to performance than optimizations at the thread level since accessing device and shared memory is orders of magnitude more expensive than accessing register file, Mirage’s $\mu$Graph generator employs a hybrid approach, considering all possible graphs up to a certain size at the kernel and block levels, and using a rule-based strategy to construct graphs at the thread level, which reduces the search space while retaining most performance optimizations. Second, to further prune the search space, Mirage introduces a pruning technique based on an abstraction of $\mu$Graphs called abstract expression, which reduces the number of $\mu$Graphs Mirage has to consider while providing a theoretical guarantee on the optimality of the discovered $\mu$Graphs.
4.1. Kernel and Block Graph Generation
[…]
Mirage checks tensor shape (14) and memory usage (15) before adding an operator, ensuring a valid prefix. A prefix passes the memory usage check if: (1) all tensors in the kernel graph can reside in device memory; and (2) all tensors in each block graph can fit in shared memory.
@@ 4.1. Kernel and Block Graph Generation
4.2. Thread Graph Construction
Mirage constructs thread graphs in a way similar to operator fusion, as shown in the third part of Figure 5 and 8-12 in Algorithm 1. Specifically, Mirage has a set of pre-defined patterns $\{(G_{i},O_{i})\}$ where each $G_{i}$ is a graph consisting of block operators and $O_{i}$ is a thread graph. Given a $\mu$Graph, Mirage traverses all its block graphs, replacing any subgraphs that match $G_{i}$ with an operator defined by $O_{i}$.
Table 1. Operators supported by Mirage. The second column shows the levels of graphs supporting the operator (K, B and T stand for kernel, block and thread graphs, respectively). The last column defines the abstract expressions of the outputs of each operator. $E$ is the mapping from tensors to the corresponding abstract expressions.
@@ 4.2. Thread Graph Construction
[…]
@@
4.3. Pruning via Abstract Expressions
When searching the space of possible $\mu$Graphs, we aim to avoid $\mu$Graph prefixes whose intermediate results cannot contribute to the desired computation. For example, for the input program $X\cdot Z+Y\cdot Z$, we can prune a prefix that computes $X\cdot Y$, but we should not prune one that computes $X+Y$, as $(X+Y)\cdot Z$ is equivalent to the input program. However, how can we determine whether a prefix can contribute to a desired computation while searching for that computation? Below, we develop a pruning technique driven by this intuition that circumvents the “chicken and egg” problem via abstraction. First, we present the abstraction — abstract expressions — and then explain how it is used for pruning. Finally, we offer a theoretical guarantee that under certain conditions, this pruning does not exclude the optimal $\mu$Graph.
Figure 6. Illustration of abstract expressions. The abstract expressions of tensors are annotated on edges. A human-friendly notation is used here: $\mathrm{e}^{a}$ denotes $\mathsf{exp}(a)$, $\sum_{k}a$ denotes $\mathsf{sum}(k,a)$, $a/b$ denotes $\mathsf{div}(a,b)$, and $a*b$ denotes $\mathsf{mul}(a,b)$. The tensors $I_{1}$, $I_{2}$ and $O$ are all $64\times 64$ matrices.
@@ 4.3. Pruning via Abstract Expressions
Recall that each edge in a $\mu$Graph corresponds to a tensor-valued function of the input tensors. Intuitively, abstract expressions abstract these functions by ignoring the differences between elements from the same input tensor. Formally, abstract expressions are first-order logic terms over the theory of integers and uninterpreted functions. In a $\mu$Graph, the abstract expression of each edge, denoted by $\mathrm{E}(\cdot)$, is defined as shown in Table 1. During the computation of a $\mu$Graph’s abstract expression, all graph-defined operators are “inlined”. Specifically, the expressions computed for a graph-defined operator’s inputs serve as inputs to the lower-level graph, and the expressions for the outputs of this lower-level graph become the output expressions of the graph-defined operator. Figure 6 shows the abstract expressions for a subgraph of attention.
While abstract expressions capture some information about the function computed at each edge, they also abstract away much of it. For example, if $X$ is a $k\times k$ matrix, summing over the rows and summing over the columns lead to the same abstract expression—$\mathsf{sum}(k,\mathrm{E}(X))$. But keeping $k$ as part of the abstract expression is crucial for effective pruning.
@@
We use abstract expressions to prune the search space of $\mu$Graphs by formalizing two relations over abstract expressions: equivalence and abstract subexpression. We prune every $\mu$Graph prefix whose abstract expression is not a subexpression of some abstract expression equivalent to that of the input program. We formalize abstract expressions as uninterpreted functions in first-order logic over the theory of integer arithmetic and uninterpreted functions, and use an SMT solver to reason about them based on two sets of axioms in Table 2: $A_{\text{eq}}$ and $A_{\text{sub}}$. $A_{\text{eq}}$ axiomatizes equivalence between abstract expressions. As will become clear below, these axioms need not be sound—it is not necessary that $\mu$Graphs whose abstract expressions are equivalent are actually equivalent. (As mentioned earlier, non-equivalent $\mu$Graphs can have identical abstract expressions.) $A_{\text{sub}}$ axiomatizes the subexpression relation between abstract expressions. The key property of this relation is that whenever a $\mu$Graph $G_{1}$ is a prefix of $G_{2}$, meaning $G_{2}$ can be obtained from $G_{1}$ by adding more operators, then $\mathrm{E}(G_{1})$ is an abstract subexpression of $\mathrm{E}(G_{2})$, i.e., $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G_{1}),\mathrm{E}(G_{2}))$, where $\models$ denotes entailment modulo the theory of integer arithmetic and uninterpreted functions.
@@
During the search, Algorithm 1 first computes the abstract expression of the input Lax program, $E_{O}$, and prunes away any $\mu$Graph prefix $G$ if $A_{\text{eq}}\cup A_{\text{sub}}\not\models\mathsf{subexpr}(\mathrm{E}(G),E_{O})$. That is, we prune a graph if its abstract expression is not a subexpression of $E_{O}$. The check is performed by invoking an SMT solver (Z3 (z3)). As an optimization, check results are cached and reused, since during the search Mirage may encounter multiple $\mu$Graphs with identical abstract expressions and SMT queries are relatively expensive.
@@
Table 2. Axiomatization of abstract expressions used for pruning. Mirage checks whether an abstract expression $E_{1}$ is a subexpression of $E_{2}$ by querying an SMT solver to check if $\mathsf{subexpr}(E_{1},E_{2})$ is entailed by these axioms. All variables are universally quantified.
Abstract Expression Property
Equivalence Axioms $A_{\text{eq}}$ Subexpression Axioms $A_{\text{sub}}$
$\mathsf{add}(x,y)=\mathsf{add}(y,x)$ $\mathsf{subexpr}(x,\mathsf{add}(x,y))$
$\mathsf{mul}(x,y)=\mathsf{mul}(y,x)$ $\mathsf{subexpr}(x,\mathsf{mul}(x,y))$
$\mathsf{add}(x,\mathsf{add}(y,z))=\mathsf{add}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{div}(x,y))$
$\mathsf{mul}(x,\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{mul}(x,y),z)$ $\mathsf{subexpr}(y,\mathsf{div}(x,y))$
$\mathsf{add}(\mathsf{mul}(x,z),\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{exp}(x))$
$\mathsf{add}(\mathsf{div}(x,z),\mathsf{div}(y,z))=\mathsf{div}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{sum}(i,x))$
[…]
$x=\mathsf{sum}(1,x)$ $\mathsf{subexpr}(x,x)$
$\mathsf{sum}(i,\mathsf{sum}(j,x))=\mathsf{sum}(i*j,x)$ $\mathsf{subexpr}(x,y)\wedge\mathsf{subexpr}(y,z)$
$\mathsf{sum}(i,\mathsf{add}(x,y))=\mathsf{add}(\mathsf{sum}(i,x),\mathsf{sum}(i,y))$ $\ \to\mathsf{subexpr}(x,z)$
$\mathsf{sum}(i,\mathsf{mul}(x,y))=\mathsf{mul}(\mathsf{sum}(i,x),y)$
$\mathsf{sum}(i,\mathsf{div}(x,y))=\mathsf{div}(\mathsf{sum}(i,x),y)$
@@
By Tables 1 and 2, we show that for any operator op, if $Y=\mbox{\rm\em op}(X_{1},\ldots,X_{n})$ then $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(X_{i}),\mathrm{E}(Y))$ (for $1\leq i\leq n)$. That is, an input to an operator is always an abstract subexpression of its output. By the reflexivity and transitivity axioms included in $A_{\text{sub}}$, it follows for any $G^{\prime}$ that is a prefix of $G$, $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G))$. Together with the assumption that $A_{\text{eq}}\models E(G_{0})=E(G)$, it follows that $A_{\text{eq}}\cup A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G_{0}))$. Thus, any prefix of $G$ will not be pruned, and $G$ will be generated by Mirage. ∎
@@
The theorem highlights the role of abstract expressions in solving the “chicken and egg” problem outlined above. To decide if a prefix $\mu$Graph is useful, we reason about whether it is a prefix of a useful computation in the abstract. The choice of the abstraction and of the axioms $A_{\text{eq}}$ represents a tradeoff between optimality and pruning. As Theorem 1 shows, we are only guaranteed to find the optimal $\mu$Graph such that $A_{\text{eq}}$ imply equivalence of its abstract expression to that of the input program. The stronger the axioms, the theorem covers more $\mu$Graphs, but we also get less pruning because more prefixes would pass the subexpression test. In particular, note that $A_{\text{eq}}$ does not consider cancellation (e.g., $\mathsf{div}(\mathsf{mul}(x,y),y)=y$). As a result, Mirage may miss some equivalent $\mu$Graphs. But including an axiom for cancellation of division and multiplication would make everything a subexpression of everything, therefore nulling the desired pruning. As our evaluation shows, our choice of $A_{\text{eq}}$ yields a good balance between pruning and optimality.
5. Probabilistic Equivalence Verifier

### v3

@@ 3 Case Study: RMSNorm
Figure 5: An overview of the $\mu$Graph generator.
RMSNorm is often followed by a matrix multiplication (MatMul). Figure 3a shows the computation graph of an RMSNorm followed by a MatMul operator, where $X$ is the input tensor, and $G$ and $W$ denote two weight tensors. Existing ML compilers generally launch two separate kernels for RMSNorm and MatMul computations, since both operations internally perform reductions across an input dimension, making it challenging to fuse their computations into a single kernel. This approach requires storing intermediate results (i.e., $Y$) in device memory since different kernels cannot share data in shared memory or register files.
@@
We highlight the key differences between the $\mu$Graph discovered by Mirage and the original $\mu$Graph. These differences involve discovering new custom kernels and combining algebraic and schedule transformations, making it infeasible to discover the final $\mu$Graph by separately considering algebraic and schedule transformations. First, Mirage reorders MatMul and the division of RMSNorm by leveraging the commutativity of matrix multiplication and element-wise division (algebraic transformation). Second, Mirage performs the accumulation in the root mean square (i.e., $A_{i}=\sum_{j}X_{ij}^{2}$) and the accumulation in the matrix multiplication (i.e., $B_{ik}=\sum_{j}X_{ij}G_{j}W_{jk}$) in parallel (schedule transformation), avoiding writing the accumulation results to device memory. Next, Mirage instantiates a thread graph to perform a sequence of element-wise operators while maintaining all intermediate results in register files (schedule transformation). Finally, the best discovered $\mu$Graph uses a new custom kernel to fuse the computation of RMSNorm and MatMul, reducing device memory access and kernel launch overheads. This $\mu$Graph outperforms the hand-written kernels in existing systems by 1.5$\times$ and 1.9$\times$ on NVIDIA A100 and H100 GPUs respectively.
4 Expression-Guided $\mu$Graph Generator
@@
3:$E_{O}\leftarrow E(G_{\mathsf{ref}})$
4:$\mathcal{S}_{0},\mathcal{S}\leftarrow\varnothing$
5:GenerateNextKernelOperator($\mathrm{Inputs}(G_{\mathsf{ref}})$)
6:for all $G\in\mathcal{S}_{0}$ do
7:  $\mathcal{S}\leftarrow\mathcal{S}\cup\{\textsc{ThreadGraphConstruction}(G)\}$
8:
9:function GenerateNextKernelOperator($G_{\mathsf{K}}$)
10:  $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{G_{\mathsf{K}}\}$
11:  for all kernel graph op type $t$; input set $I$ do
[…]
13:    if $t$ is a pre-defined operator then
14:     if $o:=\textsc{ConstructOp}(G_{\mathsf{K}},I,t)$ is valid then
15:      GenerateNextKernelOperator($G_{\mathsf{K}}\cup\{o\}$)
16:    $\triangleright$ $t$ is a graph-defined operator
@@
19:      GenerateNextBlockOperator($G_{\mathsf{K}},G_{\mathsf{B}}$)
20:
21:function GenerateNextBlockOperator($G_{\mathsf{K}},G_{\mathsf{B}}$)
22:  if all shared tensors in $G_{\mathsf{B}}$ are consumed then
23:   if $o:=\textsc{ConstructOp}(G_{\mathsf{K}},G_{\mathsf{B}}.I,G_{\mathsf{B}})$ is valid then
24:    GenerateNextKernelOperator($G_{\mathsf{K}}\cup\{o\}$)
25:  for all block graph op type $t$; input set $I$ do
[…]
27:    if $o:=\textsc{ConstructOp}(G_{\mathsf{B}},I,t)$ is valid then
28:     GenerateNextBlockOperator($G_{\mathsf{K}},G_{\mathsf{B}}\cup\{o\}$)
29:
@@
32:  if $\textsc{Subexpr}(E,E_{O})$ then $\triangleright$ Prune via abstract expressions
@@
34:   if $S.\mathrm{valid}$, $G_{\mathsf{}}.\mathrm{mAlloc}+S.\mathrm{size}\leq G_{\mathsf{}}.\mathrm{mLimit}$ then $\triangleright$ Check memory
@@
36:  return Invalid
37:
38:function ThreadGraphConstruction($G$)
39:  $G_{\mathsf{fused}}\leftarrow G$
40:  while $\exists o\in G_{\mathsf{fused}}$ that can be fused with a preceding operator do
41:   $G_{\mathsf{fused}}\leftarrow\textsc{FuseOp}(G_{\mathsf{fused}},o)$
42:  return $G_{\mathsf{fused}}$
This section introduces the Mirage $\mu$Graph generator, which automatically discovers potential $\mu$Graphs for an input tensor program. To generate $\mu$Graphs that capture optimizations at the kernel, block, and thread levels, Mirage must explore a significantly larger search space than existing superoptimizers, which only consider optimizations at the kernel level. Mirage employs two key techniques to address this challenge. First, based on the observation that optimizations at the kernel and block levels are substantially more critical to performance than optimizations at the thread level—since accessing device and shared memory is orders of magnitude more expensive than accessing register file—Mirage’s $\mu$Graph generator employs a hybrid approach: it exhaustively considers all possible graphs up to a certain size at the kernel and block levels, and uses a rule-based strategy to construct graphs at the thread level. This approach reduces the search space while retaining most performance-critical optimizations. Second, to further prune the search space, Mirage introduces a pruning technique based on an abstraction of $\mu$Graphs called abstract expression, which reduces the number of $\mu$Graphs Mirage must consider while providing a certain theoretical guarantee on the optimality of the discovered $\mu$Graphs. We introduce the hybrid $\mu$Graph generation algorithm in §4.1 and §4.2, and the expression-guided pruning techniques in §4.3.
4.1 Kernel and Block Graph Generation
Mirage generates kernel and block graphs incrementally and leverages several pruning techniques to reduce the search space, as shown in the second part of Figure 5. Specifically, Mirage maintains a prefix of a valid $\mu$Graph and iteratively extends it with new operators. For a graph $G=(V,E)$ we say that $G^{\prime}=(V^{\prime},E^{\prime})$ is a prefix of $G$ if it is a subgraph of $G$ such that $\forall u\in V^{\prime},\forall(v,u)\in E,v\in V^{\prime}$.
To generate the next operator in the kernel graph, Mirage enumerates the kernel operator type $t$ and the input tensor set $I$. If $t$ represents the graph-defined operator type, Mirage generates the associated block graph that defines its kernel computation by (1) enumerating the grid and for-loop dimensions (introduced in §2), which enables Mirage to calculate the input tensor shapes of the block graph; and (2) performing a nested generation procedure similar to that used at the kernel level but without considering graph-defined operators. 9-19 and 21-28 in Algorithm 1 show how Mirage generates kernel and block operators, respectively. Mirage checks tensor shape (33) and memory usage (34) before adding an operator, ensuring a valid prefix.
@@ 4.1 Kernel and Block Graph Generation
4.2 Thread Graph Construction
While a similar nested generation strategy can be applied to thread graphs, Mirage instead constructs them using a transformation-based approach (see the third panel of Figure 5 and lines 67 in Algorithm 1) to reduce the search space. Mirage applies operator fusion when constructing thread graphs, which reduces access to shared memory by reusing tensors in register file whenever possible. For example, Mirage fuses the three element-wise operators (Mul, Sqrt, and Div) in Figure 3b into a thread graph, avoiding saving intermediate results to shared memory and keeping the entire computation of these operators in register file. While our current implementation focuses on operator fusion, additional rule-based transformations can be used to construct thread graphs.
Table 1: Operators supported by Mirage. The second column shows the graph levels supporting each operator (K, B and T denote kernel, block, and thread graphs, respectively). The last column defines the abstract expressions for the outputs of each operator, where $E$ maps tensors to their abstract expressions.
@@ 4.2 Thread Graph Construction
Sqrt K, B $\mathrm{E}(\mathrm{Sqrt}(X))=\mathsf{sqrt}(\mathrm{E}(X))$
SiLU K, B $\mathrm{E}(\mathrm{SiLU}(X))=\mathsf{silu}(\mathrm{E}(X))$
@@
4.3 Pruning via Abstract Expressions
When searching the space of possible $\mu$Graphs, we aim to avoid $\mu$Graph prefixes whose intermediate results cannot contribute to the desired computation. For example, for the input program $X\cdot Z+Y\cdot Z$, we can prune a prefix that computes $X\cdot Y$, but we should not prune one that computes $X+Y$, as $(X+Y)\cdot Z$ is equivalent to the input program. However, how can we determine whether a prefix can contribute to a desired computation while searching for that computation? Below, we develop a pruning technique driven by this intuition that circumvents the “chicken and egg” problem via abstraction. We first present the abstraction—abstract expressions—and then explain how it is used for pruning. Finally, we offer a theoretical guarantee that, under certain conditions, this pruning does not exclude the optimal $\mu$Graph.
@@ 4.3 Pruning via Abstract Expressions
Recall that an edge in a $\mu$Graph corresponds to a tensor-valued function of the input tensors. Intuitively, abstract expressions abstract these functions by ignoring the differences between elements of the same input tensor. Formally, abstract expressions are first-order logic terms over the theory of integers and uninterpreted functions. In a $\mu$Graph, the abstract expression of each edge, denoted by $\mathrm{E}(\cdot)$, is defined in Table 1. When computing a $\mu$Graph’s abstract expression, all graph-defined operators are “inlined”. Specifically, the expressions computed for a graph-defined operator’s inputs are passed into its lower-level graph, and the resulting output expressions of that lower-level graph become the output expressions of the graph-defined operator. Figure 6 shows the abstract expressions for a subgraph of attention.
While abstract expressions capture some information about the function computed at each edge, they also abstract away many details. For example, if $X$ is a $k\times k$ matrix, summing over the rows and summing over the columns both yield the same abstract expression—$\mathsf{sum}(k,\mathrm{E}(X))$. But keeping $k$ as part of the abstract expression is crucial for effective pruning.
Figure 6: Illustration of abstract expressions. The abstract expressions of tensors are annotated on edges. A human-friendly notation is used here: $\mathrm{e}^{a}$ denotes $\mathsf{exp}(a)$, $\sum_{k}a$ denotes $\mathsf{sum}(k,a)$, $a/b$ denotes $\mathsf{div}(a,b)$, and $a*b$ denotes $\mathsf{mul}(a,b)$. The tensors $I_{1}$, $I_{2}$ and $O$ are all $64\times 64$ matrices.
@@
We use abstract expressions to prune the search space of $\mu$Graphs by formalizing two relations over abstract expressions: equivalence and abstract subexpression. Specifically, we prune any $\mu$Graph prefix whose abstract expression is not a subexpression of some abstract expression equivalent to that of the input program. We formalize abstract expressions as uninterpreted functions in first-order logic over the theory of integer arithmetic and uninterpreted functions, and use an SMT solver to reason about them based on two sets of axioms in Table 2: $A_{\text{eq}}$ and $A_{\text{sub}}$.
First, $A_{\text{eq}}$ axiomatizes equivalence between abstract expressions. As will become clear below, these axioms need not be sound—it is not required that $\mu$Graphs with equivalent abstract expressions are functionally equivalent, since non-equivalent $\mu$Graphs can have the same abstract expression. Second, $A_{\text{sub}}$ axiomatizes the subexpression relation between abstract expressions. A key property of $A_{\text{sub}}$ is that whenever a $\mu$Graph $G_{1}$ is a prefix of $G_{2}$—meaning $G_{2}$ can be constructed by extending $G_{1}$ with additional operators—$\mathrm{E}(G_{1})$ is an abstract subexpression of $\mathrm{E}(G_{2})$; formally, $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G_{1}),\mathrm{E}(G_{2}))$, where $\models$ denotes entailment modulo the theory of integer arithmetic and uninterpreted functions.
@@
During the search, Algorithm 1 first computes the abstract expression of the input Lax program, denoted $E_{O}$, and prunes any $\mu$Graph prefix $G$ if $A_{\text{eq}}\cup A_{\text{sub}}\not\models\mathsf{subexpr}(\mathrm{E}(G),E_{O})$. That is, a graph is pruned if its abstract expression is not a subexpression of $E_{O}$. This check is performed using an SMT solver (Z3 [18]). As an optimization, the results of these checks are cached and reused, since Mirage may encounter multiple $\mu$Graphs with identical abstract expressions during the search.
@@
Table 2: Axiomatization of abstract expressions used for pruning. Mirage checks whether an abstract expression $E_{1}$ is a subexpression of $E_{2}$ by querying an SMT solver to check if $\mathsf{subexpr}(E_{1},E_{2})$ is entailed by these axioms. All variables in these axioms are universally quantified.
Abstract Expression Property Comment
Equivalence Axioms $A_{\text{eq}}$
$\forall x,y.\ \mathsf{add}(x,y)=\mathsf{add}(y,x)$ commutativity
$\forall x,y.\ \mathsf{mul}(x,y)=\mathsf{mul}(y,x)$ commutativity
$\forall x,y,z.\ \mathsf{add}(x,\mathsf{add}(y,z))=\mathsf{add}(\mathsf{add}(x,y),z)$ associativity
$\forall x,y,z.\ \mathsf{mul}(x,\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{mul}(x,y),z)$ associativity
$\forall x,y,z.\ \mathsf{add}(\mathsf{mul}(x,z),\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{add}(x,y),z)$ distributivity
$\forall x,y,z.\ \mathsf{add}(\mathsf{div}(x,z),\mathsf{div}(y,z))=\mathsf{div}(\mathsf{add}(x,y),z)$ associativity
$\forall x,y,z.\ \mathsf{mul}(x,\mathsf{div}(y,z))=\mathsf{div}(\mathsf{mul}(x,y),z)$ associativity
$\forall x,y,z.\ \mathsf{div}(\mathsf{div}(x,y),z)=\mathsf{div}(x,\mathsf{mul}(y,z))$ associativity
$\forall x.\ x=\mathsf{sum}(1,x)$ identity reduction
$\forall x,i,j.\ \mathsf{sum}(i,\mathsf{sum}(j,x))=\mathsf{sum}(i*j,x)$ associativity
$\forall x,y,i.\ \mathsf{sum}(i,\mathsf{add}(x,y))=\mathsf{add}(\mathsf{sum}(i,x),\mathsf{sum}(i,y))$ associativity
$\forall x,y,i.\ \mathsf{sum}(i,\mathsf{mul}(x,y))=\mathsf{mul}(\mathsf{sum}(i,x),y)$ distributivity
$\forall x,y,i.\ \mathsf{sum}(i,\mathsf{div}(x,y))=\mathsf{div}(\mathsf{sum}(i,x),y)$ distributivity
$\forall x,y.\ \mathsf{mul}(\mathsf{exp}(x),\mathsf{exp}(y))=\mathsf{exp}(\mathsf{add}(x,y))$ distributivity
$\forall x,y.\ \mathsf{mul}(\mathsf{sqrt}(x),\mathsf{sqrt}(y))=\mathsf{sqrt}(\mathsf{mul}(x,y))$ distributivity
Subexpression Axioms $A_{\text{sub}}$
$\forall x,y.\ \mathsf{subexpr}(x,\mathsf{add}(x,y))$
$\forall x,y.\ \mathsf{subexpr}(x,\mathsf{mul}(x,y))$
$\forall x,y.\ \mathsf{subexpr}(x,\mathsf{div}(x,y))$
$\forall x,y.\ \mathsf{subexpr}(y,\mathsf{div}(x,y))$
$\forall x.\ \mathsf{subexpr}(x,\mathsf{exp}(x))$
$\forall x.\ \mathsf{subexpr}(x,\mathsf{sqrt}(x))$
$\forall x.\ \mathsf{subexpr}(x,\mathsf{silu}(x))$
$\forall x,i.\ \mathsf{subexpr}(x,\mathsf{sum}(i,x))$
$\forall x.\ \mathsf{subexpr}(x,x)$ reflexivity
$\forall x,y,z.\ \mathsf{subexpr}(x,y)\wedge\mathsf{subexpr}(y,z)\to\mathsf{subexpr}(x,z)$ transitivity
@@
By Tables 1 and 2, we show that for any operator op, if $Y=\mbox{\rm\em op}(X_{1},\ldots,X_{n})$, then $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(X_{i}),\mathrm{E}(Y))$ for $1\leq i\leq n$. That is, the abstract expression of each input to op is always a subexpression of op’s output. Given that $A_{\text{sub}}$ includes reflexivity and transitivity axioms, it follows that for any $G^{\prime}$ that is a prefix of $G$, $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G))$. Together with the assumption that $A_{\text{eq}}\models E(G_{0})=E(G)$, we have $A_{\text{eq}}\cup A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G_{0}))$. Thus, no prefix of $G$ will be pruned, and Mirage will generate $G$. ∎
@@
The theorem highlights the role of abstract expressions in solving the “chicken and egg” problem outlined above. To decide if a prefix $\mu$Graph is useful, we reason about whether it is a prefix of a useful computation in the abstract. The choice of abstraction and the axioms $A_{\text{eq}}$ represents a tradeoff between optimality and pruning. As Theorem 1 shows, we are only guaranteed to find the optimal $\mu$Graph whose abstract expression is equivalent to that of the input program under $A_{\text{eq}}$. Stronger axioms expand the set of $\mu$Graphs covered by the theorem but reduce pruning effectiveness, since more prefixes would pass the subexpression test. In particular, note that $A_{\text{eq}}$ does not include cancellation rules (e.g., $\mathsf{div}(\mathsf{mul}(x,y),y)=y$). As a result, Mirage may miss some equivalent $\mu$Graphs. However, including such axioms would make everything a subexpression of everything, therefore nulling desired pruning. As our evaluation shows, the chosen $A_{\text{eq}}$ yields a good balance between pruning and optimality.
5 Probabilistic Equivalence Verifier